OpenStudy (anonymous):

Textbook says the integral of (sin cos^3) is -(cos^4)/4 . I'm using trig integrals technique which instead gives me (sin^2)/2 - (sin^4)/4. What gives?

OpenStudy (rational):

\[\large \int \sin x \,\cos^3x \, dx\]

OpenStudy (rational):

like that ?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

what technique are you using ? could you elaborate

OpenStudy (anonymous):

Sure, I turned cos^3 into cos cos^2. Then turned cos^2 into (1-sin^2). u-subbed sin and integrated to the above answer

OpenStudy (rational):

Looks legit! so with your method, you get something like this ?\[\int\sin x~\cos x~(\cos^2x)\, dx =\int\sin x~\cos x~(1-\sin^2x)\, dx= \int u(1-u^2)\, du\]

OpenStudy (anonymous):

exactly

OpenStudy (rational):

\[ \int u(1-u^2)\, du=\frac{u^2}{2}-\frac{u^4}{4}+C = \frac{\sin^2x}{2}-\frac{\sin^4x}{4}+C\]

OpenStudy (rational):

you want to match your answer wid textbook is it ?

OpenStudy (rational):

btw, your answer is also correct

OpenStudy (anonymous):

Yea, i plugged in some angles to see if maybe it would result in the same answer but no dice. They still don't match :(

OpenStudy (rational):

they match upto a " constant "

OpenStudy (rational):

let me ask you a quick question, what are the derivatives of below two functions ? 1) \(f(x)=\sin x\) 2) \(g(x) = \sin x+100000\)

OpenStudy (anonymous):

both are cosx

OpenStudy (rational):

so whats "the" antiderivative of \(\cos x\) ?

OpenStudy (anonymous):

sinx

OpenStudy (anonymous):

sorry, i'm not sure where this is leading lol

OpenStudy (rational):

the question is actually tricky how do you know the antiderivative of \(\cos x\) is not \(\sin x+100000\) ?

OpenStudy (anonymous):

Hmm.. well i guess it seems a little counter intuitive that the functions wouldn't be equal to each other given the same angle. So then both answers are correct?

OpenStudy (rational):

Yes both answers are correct because they differ only by a constant.

OpenStudy (rational):

It would be fun to actually try and prove above statement

OpenStudy (anonymous):

Good ol' calculus. Thanks for your help :)

OpenStudy (rational):

show that "your answer" = "textbook answer" + C

OpenStudy (anonymous):

Lol it'll be fun when I don't have a final breathing down my neck.

OpenStudy (rational):

It is important to acknowledge that the antiderivative is not a single function, but a family of functions, all differing by a constant.

OpenStudy (rational):

\[\int \cos x\, dx = \sin x + C\] where \(C\) is an arbitrary constant

OpenStudy (anonymous):

Agreed - i like to know the why's. Not just the how's. Thanks for your time ! :)

OpenStudy (sdfgsdfgs):

@rational they are...cos^4 = (1-sin^2)^2 =1-2sin^2+sin^4 so -cos^4/4=1/4+sin^2/2-sin^4/4

OpenStudy (rational):

Ahh nice xD so not so hard to show they differ by a constant : 1/4

OpenStudy (anonymous):

This example illustrates a pitfall of following the textbook's solution.

OpenStudy (campbell_st):

well there seems to be a little problem in the interpretation of the question i'd do this \[u = \cos(x) ~~~~\frac{du}{dx} = -\sin(x) ~~~~so ~~~~du = -\sin(x)dx\] so the problem, rewritten is \[\int\limits \cos^3(x) \sin(x)~dx\] when when using substitution becomes \[\int\limits u^3 \times -du!!~~~or ~~~-\int\limits u^3 du\] which becomes \[-\frac{u^4}{4} + c\] make the substitution for u and you get \[-\frac{\cos^4(x)}{4}\] this is a high school maths question and solution

OpenStudy (campbell_st):

hope it makes sense... it required substitution, but just a simple substitution.

OpenStudy (campbell_st):

oops forgot to add the constant C to the solution

OpenStudy (anonymous):

@campbell_st that would be the most direct substituion

OpenStudy (anonymous):

@campbell_st it makes perfect sense! I'm beating myself up for not even considering that route...but this was overall a good exercise to see it worked out from different angles.